OpenStudy (shamil98):

sketch the curve with the given polar equation by first sketching the graph of r as a function of theta in Cartesian coordinates. r = -2 sin theta so i get how to graph it in cartesian coordinates but not how to sketch the curve thereafter..

OpenStudy (anonymous):

Is there a difference between graphing and sketching that I'm missing out on?

OpenStudy (shamil98):

same thing

OpenStudy (shamil98):

second part is asking to graph it in polar coordinates by using the one from the first part

OpenStudy (anonymous):

I'm a little inclined towards rectangular coordinates, so oh well. :P \[r^2 = -2r \sin \theta\]\[\Rightarrow x^2 + y^2 = -2x\]\[\Rightarrow x^2 + 2x + y^2 = 0\]\[\Rightarrow (x+1)^2 + y^2 = 1\]Familiar much?

OpenStudy (anonymous):

Ugh, I meant\[x^2 + (y+1)^2 = 1 \]

OpenStudy (anonymous):

Because \(r \sin \theta = y\), not \(x\). That was a bad mistake on my part.

OpenStudy (shamil98):

oh right i forgot bout that method lol thanks

OpenStudy (anonymous):

no problem but you wouldn't be thanking us when the bangladeshi n00bs lose to us in the asia cup final <3